OpenStudy (anonymous):

(Medal+Fan+Cookie ^_^) helppp Solve. 3(h – 4) = –1/2(24 – 6h)

OpenStudy (*insert name here*):

Start by distributing.

OpenStudy (anonymous):

3(h-4) = -1/2(24 - 6h) 3h - 12 = -12 + 3h -3h. -3h -12. = -12 So h = any Real number

OpenStudy (anonymous):

okay. and then what? O_O

OpenStudy (anonymous):

ooo wait that's how you do it?

OpenStudy (anonymous):

yeah